OpenStudy (anonymous):

A high-altitude spherical weather balloon expands as it rises due to the drop in atmospheric pressure. Suppose that the radius r increases at the rate of 0.02 inches per second and that r = 36 inches at time t = 0. Determine the equation that models the volume V of the balloon at time t and find the volume when t = 360 seconds.

OpenStudy (anonymous):

a. V(t) = 4π(0.02t)2; 651.44 in.3 b. V(t) = 4π(36 + 0.02t)2; 1,694,397.14 in.3 c. V(t) = ; 4690.37 in.3 d. V(t) = ; 337,706.83 in.3 These are the answer choices!:)

OpenStudy (abb0t):

First of all, do you know the volume of a sphere? \[V = \frac{ 4 }{ 3 } \pi r^3 \]

OpenStudy (abb0t):

At the time it is first being launched, time (t) equals zero.

OpenStudy (anonymous):

okay i got 195432.1958 of V....

OpenStudy (abb0t):

and rate is changing, hence you have (rt) - radius changes with time, which means you're multiplying. add this change to the radius. it adds up.

OpenStudy (abb0t):

I don't know if answer two is correct, because volume of a sphere is r cubed. and u have r squared.

OpenStudy (abb0t):

You should get: \[V = \frac{ 4 }{ 3 }\pi (r+∆rt)^3\]

OpenStudy (anonymous):

@abb0t oh so it will be the choice b right?

OpenStudy (anonymous):

Hmmm.

OpenStudy (anonymous):

Basically, I would solve for \(r(t)\) using \[ \int_0^{t}0.02du = \int_0^t\frac{dr}{du}du = r(t)-r(0) = r(t) - 36\implies r(t) = \int_0^t0.02dt+36 \]

OpenStudy (anonymous):

Clearly \[ V(t) = \frac{4}{3}[r(t)]^3 \]As @abb0t said.